The current supplied to an air conditioner unit is 6.5 amps. The air conditioner is wired using a 10-gauge (diameter 2.588 mm) wire. The charge density is n = 5.12× 1028 electrons/m3. Find the magnitude of the drift velocity.


Round your answer to 6 decimal places


1
Expert's answer
2021-05-13T09:20:39-0400

The magnitude of the drift velocity is


"v=\frac{I}{neA},\\\space\\
v=\frac{6.5}{(5.12·10^{28})(1.6·10^{-19})(\pi(2.588·10^{-3})^2/4)}=0.000151\text{ m/s}."